# Exact Gauss-Newton Optimization for Training Deep Neural Networks

### v1

@@ Abstract
We present EGN, a stochastic second-order optimization algorithm that combines the generalized Gauss-Newton (GN) Hessian approximation with low-rank linear algebra to compute the descent direction. Leveraging the Duncan-Guttman matrix identity, the parameter update is obtained by factorizing a matrix which has the size of the mini-batch. This is particularly advantageous for large-scale machine learning problems where the dimension of the neural network parameter vector is several orders of magnitude larger than the batch size. Additionally, we show how improvements such as line search, adaptive regularization, and momentum can be seamlessly added to EGN to further accelerate the algorithm. Moreover, under mild assumptions, we prove that our algorithm converges to an $\epsilon$-stationary point at a linear rate. Finally, our numerical experiments demonstrate that EGN consistently exceeds, or at most matches the generalization performance of well-tuned SGD, Adam, and SGN optimizers across various supervised and reinforcement learning tasks.
@@ 1 Introduction
Optimization plays a pivotal role in Machine Learning (ML), with gradient-based methods being at the forefront. Stochastic Gradient Descent (SGD) [61], a first-order stochastic optimization algorithm, and its accelerated versions such as momentum-based approaches [51, 67, 44, 16], adaptive learning rates [22, 30, 81], and a combination of the two [36, 21, 80], have been instrumental in numerous ML applications. For example, in Computer Vision (CV) ResNets [29] are trained with SGD, AdaGrad [22] is used for training recommendation systems [50], language models GPT-3 [12] and LLaMA [69] are optimized with Adam [36] and AdamW [43], respectively. Despite their cheap and relatively easy-to-implement updates, first order-methods (FOMs) suffer from several shortcomings. FOMs are sensitive to hyper-parameter selection, and the optimal hyper-parameter set typically does not transfer well across different problems which leads to a costly procedure of hyper-parameter tuning. Also, FOMs are slow to converge in the flat regions of the loss landscape, where the Hessian is ill-conditioned [62].
@@
Second-order methods (SOMs) incorporate the (approximate) curvature information into the update in order to effectively precondition the gradient vector. In contrast to first-order algorithms, SOMs are shown to be robust to the selection of hyper-parameters [77] and to potentially offer faster convergence [2, 6]. So far, the adoption of second-order methods for ML problems has been limited due to the complexity in calculating and storing the Hessian and the computational load of solving the linear system $\mathbf{H}\mathbf{d}=-\mathbf{g}$, where $\mathbf{H}$ is the (approximate) Hessian matrix, $\mathbf{g}$ is the gradient of the loss function, and $\mathbf{d}$ is the descent direction. Addressing these computational challenges, most approaches use a combination of Hessian approximation and an efficient algorithmic technique for solving the linear system. Common approximations to the Hessian include diagonal scaling [78, 42], the empirical Fisher matrix [60], the quasi-Newton approach [65, 14, 5], and the Gauss-Newton (GN) approximation [24, 70, 13].
@@
In this work, we follow the Gauss-Newton approach to Hessian approximation. Our innovation is in the use of an efficient exact linear algebra identity—the Duncan-Guttman (DG) formula [23, 26]—to speed up the inversion of the Hessian matrix. Compared to Hessian-free optimization (HFO) [46, 47, 37] and Inexact Gauss-Newton (iGN) [70], Exact Gauss-Newton (EGN) solves the system $\mathbf{H}\mathbf{d}=-\mathbf{g}$ exactly with the same algorithmic complexity burden. The closest approach to EGN, to the best of our knowledge, is the application of the Sherman–Morrison-Woodbury (SMW) formula [60]. As we outline below, by using a different matrix identity we are able to solve for the descent direction in fewer matrix operations than SWM, thus reducing the algorithmic complexity.
@@
  • We provide a theoretical analysis of the EGN algorithm and establish an algorithmic upper bound on the complexity of EGN in finding an $\epsilon$-stationary point.
@@ 2 Preliminaries
2.1 Notation
[…]
2.2 Problem Formulation
We adopt the Empirical Risk Minimization (ERM) framework [49] and consider the problem of finding the weights $\mathbf{w}\in{\mathbb{R}}^{d}$ of a parametric function $\Phi:{\mathbb{R}}^{m}\times{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{c}$, e.g., a neural network, such that it minimizes the empirical risk over a dataset $\mathcal{D}$ consisting of $N$ pairs $(\mathbf{y}_{i},\mathbf{x}_{i})$ where $\mathbf{x}_{i}\in{\mathbb{R}}^{m}$ is a vector of features and $\mathbf{y}_{i}\in{\mathbb{R}}^{c}$, $c\geq 1$ is a target vector. We want to solve the following optimization problem:
@@ 2.3 Gradient-based Optimization
where $\alpha_{t}>0$ is a learning rate and $\mathbf{d}_{t}$ is a descent direction obtained from the gradient. In general, we have $\mathbf{d}_{t}=-\mathbf{C}_{t}\mathbf{g}_{t}$ where $\mathbf{C}_{t}$ is a preconditioning matrix that scales, rotates, and shears the gradient of the mini-batch loss $\mathbf{g}\coloneqq\nabla_{\mathbf{w}}\mathcal{L}_{b}$. Notice that by setting $b=1$, $\mathbf{C}_{t}=\mathbf{I}_{d}$ and $\mathbf{d}_{t}=-\mathbf{g}_{t}$ we recover the incremental SGD update [61]. In practice, the preferred training algorithm is often an accelerated version of mini-batch SGD.
@@
where $\mathbf{H}_{t}\coloneqq\nabla_{\mathbf{w}}^{2}\mathcal{L}_{b}$ is the Hessian matrix of the mini-batch loss. Finding the direction $\mathbf{d}_{t}$ by solving the system (2.5) using $\mathbf{H}_{t}$ or its approximation defines the broad spectrum of second-order methods. Setting $\mathbf{C}_{t}=\mathbf{H}^{-1}_{t}$ results in Newton’s method. This method suffers from several drawbacks: (a) one needs to compute second-order derivatives with respect to $\mathbf{w}$; (b) $\mathbf{C}_{t}$ has to be positive-definite to ensure descent; (c) the linear system (2.5) must be solved, which in general scales cubically with the dimension of $\mathbf{w}$; and (d) since $\mathbf{H}_{t}$ is a noisy estimate of the true Hessian of the empirical risk $\mathcal{L}$, $\mathbf{H}^{-1}_{t}$ can result in suboptimal conditioning. These issues and the fact that $\mathbf{w}$ is of rather large dimension have made the direct application of Newton’s method practically irrelevant in ML applications. Indeed, for problems like Large Language Model (LLM) pre-training [71, 12, 69] or CV tasks [29], $\mathbf{w}$ can be extremely high-dimensional, e.g., $314\cdot 10^{9}$ parameters for Grok-1 [76] model, $7\cdot 10^{9}$ to $65\cdot 10^{9}$ parameters for LLaMA-family models [69] and $0.27\cdot 10^{6}$ to $19.4\cdot 10^{6}$ parameters for ResNets [29], which explains why accelerated first-order methods are usually preferred.
In order to make SOMs scalable for ML applications one could instead approximate the inverse Hessian, i.e., $\mathbf{C}_{t}\approx\mathbf{H}^{-1}_{t}$. Examples of such preconditioning include diagonal scaling (e.g., with Hutchinson method [78, 42]) with the idea of extracting the (approximate) diagonal elements of $\mathbf{H}_{t}$ while neglecting the off-diagonal terms; the quasi-Newton approach [14, 5] that approximates the Hessian using the information from past and current gradients; and the Gauss-Newton method [24, 60] that leverages the Jacobian of the residuals, neglecting second-order cross-derivatives, particularly suitable for loss functions structured as (2.2). In this paper, we use the Gauss-Newton Hessian approximation which is shown to provide a good approximation of the true Hessian for ML applications [62, 56].
@@ 2.4 Generalized Gauss-Newton Hessian Approximation
We consider the Generalized Gauss-Newton (GGN) Hessian approximation scheme [64, 8, 56] which is suited for both regression and multi-class classification tasks.
The GGN Hessian approximation is constructed using only first-order sensitivities (see derivation in Appendix B) to obtain
@@
where we vertically stack individual Jacobians $\mathbf{J}_{\Phi_{i}}:=\frac{\partial\Phi(\mathbf{x}_{i};\mathbf{w})}{\partial\mathbf{w}}$ for each sample in the batch $\mathcal{B}$ to form $\mathbf{J}=\left[\begin{array}[]{ccc}\mathbf{J}_{\Phi_{1}}&\dots&\mathbf{J}_{\Phi_{b}}\end{array}\right]^{\top}\in{\mathbb{R}}^{bc\times d}$ and construct a block diagonal matrix $\mathbf{Q}=\text{blkdiag}(\mathbf{Q}_{\ell_{1}},\mathbf{Q}_{\ell_{2}},\ldots,\mathbf{Q}_{\ell_{b}})\in{\mathbb{R}}^{bc\times bc}$, where $\mathbf{Q}_{\ell_{i}}=\frac{\partial^{2}\ell(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w}))}{\partial\Phi^{2}}\in{\mathbb{R}}^{c\times c}$. As pointed out in [63], approximation (2.6) is justified since the true Hessian is dominated by the term $\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}$. Using the inverse of the GGN Hessian as a preconditioner yields the following direction
@@
The Gauss-Newton step (2.7) solves issues (a) and, partially, (b), since no second-order derivatives need to be computed and the Hessian approximation is positive semi-definite by construction provided that the loss function $\ell\left(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w})\right)$ is convex. A positive-definite Hessian approximation can easily be obtained from $\mathbf{H}^{\mathrm{GN}}$, e.g., by adding to it a small constant times the identity matrix. This approach is called Levenberg-Marquardt (LM) [53] and is often used in practice, such that
@@
with $\lambda>0$. The regularizer $\lambda\mathbf{I}_{d}$ serves a dual purpose: it ensures that the approximate Hessian matrix is invertible and also it helps to avoid over-fitting [49].
@@
We can show (see Appendix B.2) that the gradient and the GN Hessian for the MSE loss read
@@
As shown in Appendix B.3, the gradient and the GN Hessian for the CE loss are
@@
where $\mathbf{J}\in{\mathbb{R}}^{bc\times d}$ is a matrix of stacked Jacobians, $\mathbf{r}\in{\mathbb{R}}^{bc}$ is a vector of (pseudo-)residuals defined as $\mathbf{r}:=\left[\begin{array}[]{ccc}\left(\sigma\left(\Phi(\mathbf{x}_{1};\mathbf{w})\right)-\mathbf{y}_{1}\right)^{\top}&\dots&\left(\sigma\left(\Phi(\mathbf{x}_{b};\mathbf{w})\right)-\mathbf{y}_{b}\right)^{\top}\end{array}\right]^{\top}$, and $\mathbf{Q}\in{\mathbb{R}}^{bc\times bc}$ is a block diagonal matrix of stacked matrices $\mathbf{Q}_{\ell_{i}}$ that each have $\sigma(\mathbf{z}_{i,k})(1-\sigma(\mathbf{z}_{i,k}))$ across the diagonal and $-\sigma(\mathbf{z}_{i,k})\sigma(\mathbf{z}_{i,l})$ off-diagonal.

### v2

@@ Abstract
We present Exact Gauss-Newton (EGN), a stochastic second-order optimization algorithm that combines the generalized Gauss-Newton (GN) Hessian approximation with low-rank linear algebra to compute the descent direction. Leveraging the Duncan-Guttman matrix identity, the parameter update is obtained by factorizing a matrix which has the size of the mini-batch. This is particularly advantageous for large-scale machine learning problems where the dimension of the neural network parameter vector is several orders of magnitude larger than the batch size. Additionally, we show how improvements such as line search, adaptive regularization, and momentum can be seamlessly added to EGN to further accelerate the algorithm. Moreover, under mild assumptions, we prove that our algorithm converges in expectation to a stationary point of the objective. Finally, our numerical experiments demonstrate that EGN consistently exceeds, or at most matches the generalization performance of well-tuned SGD, Adam, GAF, SQN, and SGN optimizers across various supervised and reinforcement learning tasks.
@@ 1 Introduction
Optimization plays a pivotal role in Machine Learning (ML), with gradient-based methods being at the forefront. Stochastic Gradient Descent (SGD) robbins1951stochastic , a first-order stochastic optimization algorithm, and its accelerated versions such as momentum-based approaches nesterov1983method ; sutskever2013importance ; lucas2018aggregated ; chen2022demon , adaptive learning rates duchi2011adaptive ; hinton2012neural ; zeiler2012adadelta , and a combination of the two kingma2014adam ; dozat2016incorporating ; zaheer2018adaptive , have been instrumental in numerous ML applications. For example, in Computer Vision (CV) ResNets he2016deep are trained with SGD, AdaGrad duchi2011adaptive is used for training recommendation systems naumov2019deep , language models GPT-3 brown2020language and LLaMA touvron2023llama are optimized with Adam kingma2014adam and AdamW loshchilov2017decoupled , respectively. Despite their cheap and relatively easy-to-implement updates, first order-methods (FOMs) suffer from several shortcomings. FOMs are sensitive to hyper-parameter selection, and the optimal hyper-parameter set typically does not transfer well across different problems which leads to a costly procedure of hyper-parameter tuning. Also, FOMs are slow to converge in the flat regions of the loss landscape, where the Hessian is ill-conditioned sagun2017empirical .
@@
Second-order methods (SOMs) incorporate the (approximate) curvature information into the update in order to effectively precondition the gradient vector. In contrast to first-order algorithms, SOMs are shown to be robust to the selection of hyper-parameters xu2020second and to potentially offer faster convergence agarwal2017second ; bollapragada2019exact . So far, the adoption of second-order methods for ML problems has been limited due to the complexity of calculating and storing the Hessian and the computational load of solving the linear system $\mathbf{H}\mathbf{d}=-\mathbf{g}$, where $\mathbf{H}$ is the (approximate) Hessian matrix, $\mathbf{g}$ is the gradient of the loss function, and $\mathbf{d}$ is the descent direction. Addressing these computational challenges, most approaches use a combination of Hessian approximation and an efficient algorithmic technique for solving the linear system. Common approximations to the Hessian include diagonal scaling yao2021adahessian ; liu2023sophia , the empirical Fisher matrix ren2019efficient , the quasi-Newton approach schraudolph2007stochastic ; byrd2016stochastic ; berahas2016multi , and the Gauss-Newton (GN) approximation gargiani2020promise ; tran2020stochastic ; brust2021nonlinear .
@@
In this work, we follow the Gauss-Newton approach to Hessian approximation. We apply a special derivation inspired by adeoye2023score ; adeoye2021sc which uses an efficient exact linear algebra identity—the Duncan-Guttman (DG) formula duncan1944lxxviii ; guttman1946enlargement —to speed up the inversion of the Hessian matrix. Compared to Hessian-free optimization (HFO) martens2010deep ; martens2011learning ; kiros2013training and Inexact Gauss-Newton (iGN) tran2020stochastic , this approach allows Exact Gauss-Newton (EGN) to solve the system $\mathbf{H}\mathbf{d}=-\mathbf{g}$ exactly with the same algorithmic complexity burden. Moreover, compared to methods that directly apply the Sherman-Morrison-Woodbury (SMW) formula (see, e.g., ren2019efficient ), we solve for the descent direction in fewer matrix operations, thus reducing the algorithmic complexity.
@@
  • We provide a theoretical analysis of the EGN algorithm and establish that EGN finds a stationary point in expectation for a large enough iteration count.
@@
2 Related Work
Our method can be viewed within the broader context of approximate second-order stochastic optimization. Some notable approaches include diagonal scaling yao2021adahessian ; liu2023sophia , Krylov subspace descent vinyals2012krylov , Hessian-free optimization martens2010deep ; martens2011learning ; kiros2013training , quasi-Newton approaches schraudolph2007stochastic ; byrd2016stochastic ; berahas2016multi ; wills2021stochastic ; liu2022quasi , Gauss-Newton  botev2017practical ; gargiani2020promise and Natural Gradient amari1998natural ; kunstner2019limitations methods. A detailed overview of second-order optimization methods for large-scale machine learning problems can be found in bottou2018optimization ; sun2019survey ; xu2020second .
SGN gargiani2020promise Exact via reverse mode autodiff Approximate with CG -
LM pooladzandi2022improving Exact via reverse mode autodiff Approximate with CG Line search, momentum, uphill step acceptance
SGN tran2020stochastic Exact via reverse mode autodiff Approximate with ADPG -
SGN2 tran2020stochastic Approximate with SARAH estimators Approximate with ADPG -
NLLS1, NLLSL brust2021nonlinear Rank-1, Rank-L approximation Exact with SMW formula -
SMW-GN ren2019efficient Exact via reverse mode autodiff Exact with SMW formula Adaptive regularization
EGN (this paper) Any Jacobian estimation algorithm (exact via backpropagation is the default) Exact with DG identity (see Theorem 4.1 and Lemma 4.2) Line search, adaptive regularization, momentum
Most closely related to our work are the algorithms inspired by the Gauss-Newton approach. Such methods approximate the Hessian of the loss function using only first-order sensitivities. In practice, the damped version of the Gauss-Newton direction is often calculated, forming the stochastic Levenberg-Marquardt (SLM) group of algorithms. We can classify SLM methods by three dimensions: (a) by the type of the Jacobian estimation algorithm used; (b) by the matrix inversion algorithm; and (c) by additional adaptive parameters and acceleration techniques. Based on this paradigm we summarize selected SLM algorithms in Table 1.
The Jacobian can either be calculated exactly through the reverse mode of automatic differentiation as proposed by, e.g., gargiani2020promise ; brust2021nonlinear ; tran2020stochastic or be estimated approximately. Low rank Jacobian estimation is suggested by the NLLS1 and NLLSL algorithms brust2021nonlinear with experimental results showing almost on par performance with the exact Jacobian version of the methods. SGN2 tran2020stochastic uses SARAH estimators for approximating function values and Jacobians with SGN2 performing better than SGN tran2020stochastic that assumes the exact Jacobian. In liu2023sophia the Gauss-Newton-Bartlett (GNB) estimator is introduced to adapt the GN method to large-scale classification problems. EGN does not mandate a specific computation technique for the Jacobian. In our experiments we rely on backpropagation deferring other methods to further research.
Solving $\mathbf{H}\mathbf{d}=-\mathbf{g}$ naively for a neural network with $d$ parameters has complexity $\mathcal{O}\left(d^{3}\right)$. The procedure becomes practically infeasible even for networks of moderate size, so several alternative approaches have been proposed. Following bollapragada2019exact , we distinguish between inexact and exact solutions to the linear system. Inexact methods rely on iterative algorithms to solve the system approximately in as few iterations as possible. Among such algorithms we mention the Conjugate Gradient (CG) method used in Hessian-free optimization martens2011learning ; kiros2013training as well as in GN methods like SGN gargiani2020promise and LM pooladzandi2022improving ; the Accelerated Dual Proximal-Gradient (ADPG) method proposed in tran2020stochastic for SGN and SGN2 solvers; the Stochastic Gradient Iteration approach (SGI), analysed in bollapragada2019exact as part of the Newton-SGI solver. Exact methods are typically based on linear algebra identities. For example, in ren2019efficient and brust2021nonlinear the system is solved exactly with the Sherman-Morrison-Woodbury (SMW) formula. Contrarily, we follow  adeoye2023score and derive the EGN update formula using the Duncan-Guttman matrix identity duncan1944lxxviii ; guttman1946enlargement .
State-of-the-art implementations of the GN algorithm often include additional improvements to address the issues of stochasticity of the Hessian matrix, computational load of calculating the Jacobian and adaptive hyper-parameter tuning. Just as with the gradient, the stochastic sampling introduces noise in the Hessian which leads to the erroneous descent direction. A common solution to combat noisy estimates is to add temporal averaging (or momentum), e.g., with exponential moving averages (EMA). Examples of such approach are AdaHessian yao2021adahessian and Sophia liu2023sophia that keep EMA of a diagonal Hessian, as well as  kiros2013training that incorporates momentum into the HFO framework. The idea of re-using the Hessian estimate from previous iterations is formalized in doikov2023second showing that evaluating the Hessian “lazily” once per $k$ iterations significantly reduces the computational burden while at the same time does not degrade the performance that much. Although second-order methods typically require less tuning botev2017practical ; gargiani2020promise ; xu2020second , the SLM approach still requires setting a learning rate $\alpha$ and the regularization parameter $\lambda$. The line search for $\alpha$, common in the deterministic optimization, is problematic in the stochastic setting due to the high variance of the loss gradient norm curtis2020adaptive . Still, there are promising attempts to incorporate line search into quasi-Newton methods wills2021stochastic , HFO kiros2013training as well as Gauss-Newton pooladzandi2022improving . Adaptive regularization in a manner similar to the deterministic Levenberg-Marquardt approach is proposed in hong2020stochastic ; ren2019efficient .
@@ 3 Preliminaries
We use boldface letters to denote vectors and matrices. The $n\times n$ identity matrix is denoted by $\mathbf{I}_{n}$, and we omit the subscript when the size is clear from the context. The subscript $t$ represents the iteration within the optimization loop, and may be omitted to avoid overloading the notation. We define the standard inner product between two vectors $\mathbf{x},\mathbf{y}$ as $\langle\mathbf{x},\mathbf{y}\rangle\coloneqq\mathbf{x}^{\top}\mathbf{y}$. The standard Euclidean norm is denoted by $\|\cdot\|$, and the expected value of a random variable is denoted by $\mathbb{E}$$\cdot$].
We adopt the Empirical Risk Minimization (ERM) framework pml1Book and consider the problem of finding the weights $\mathbf{w}\in{\mathbb{R}}^{d}$ of a parametric function $\Phi:{\mathbb{R}}^{m}\times{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{c}$, e.g., a neural network, such that it minimizes the empirical risk over a dataset $\mathcal{D}$ consisting of $N$ pairs $(\mathbf{y}_{i},\mathbf{x}_{i})$ where $\mathbf{x}_{i}\in{\mathbb{R}}^{m}$ is a vector of features and $\mathbf{y}_{i}\in{\mathbb{R}}^{c}$, $c\geq 1$ is a target vector. We want to solve the following optimization problem:
@@ 3.1 Gradient-based Optimization
where $\alpha_{t}>0$ is a learning rate and $\mathbf{d}_{t}$ is a descent direction obtained from the gradient. In general, we have $\mathbf{d}_{t}=-\mathbf{C}_{t}\mathbf{g}_{t}$ where $\mathbf{C}_{t}$ is a preconditioning matrix that scales, rotates, and shears the gradient of the mini-batch loss $\mathbf{g}\coloneqq\nabla_{\mathbf{w}}\mathcal{L}_{b}$. Notice that by setting $b=1$, $\mathbf{C}_{t}=\mathbf{I}_{d}$ and $\mathbf{d}_{t}=-\mathbf{g}_{t}$ we recover the incremental SGD update robbins1951stochastic . In practice, the preferred training algorithm is often an accelerated version of mini-batch SGD.
@@
where $\mathbf{H}_{t}\coloneqq\nabla_{\mathbf{w}}^{2}\mathcal{L}_{b}$ is the Hessian matrix of the mini-batch loss. Finding the direction $\mathbf{d}_{t}$ by solving the system (3.5) using $\mathbf{H}_{t}$ or its approximation defines the broad spectrum of second-order methods. Setting $\mathbf{C}_{t}=\mathbf{H}^{-1}_{t}$ results in Newton’s method. This method suffers from several drawbacks: (a) one needs to compute second-order derivatives with respect to $\mathbf{w}$; (b) $\mathbf{C}_{t}$ has to be positive-definite to ensure descent; (c) the linear system (3.5) must be solved, which in general scales cubically with the dimension of $\mathbf{w}$; and (d) since $\mathbf{H}_{t}$ is a noisy estimate of the true Hessian of the empirical risk $\mathcal{L}$, $\mathbf{H}^{-1}_{t}$ can result in suboptimal conditioning. These issues and the fact that $\mathbf{w}$ is of rather large dimension have made the direct application of Newton’s method practically irrelevant in ML applications. Indeed, for problems like Large Language Model (LLM) pre-training vaswani2017attention ; brown2020language ; touvron2023llama or CV tasks he2016deep , $\mathbf{w}$ can be extremely high-dimensional, e.g., $314\cdot 10^{9}$ parameters for Grok-1 grok1 model, $8\cdot 10^{9}$ to $405\cdot 10^{9}$ parameters for LLaMA-family models dubey2024llama and $0.27\cdot 10^{6}$ to $19.4\cdot 10^{6}$ parameters for ResNets he2016deep , which explains why accelerated first-order methods are usually preferred.
In order to make SOMs scalable for ML applications one could instead approximate the inverse Hessian, i.e., $\mathbf{C}_{t}\approx\mathbf{H}^{-1}_{t}$. Examples of such preconditioning include diagonal scaling (e.g., with Hutchinson method yao2021adahessian ; liu2023sophia ) with the idea of extracting the (approximate) diagonal elements of $\mathbf{H}_{t}$ while neglecting the off-diagonal terms; the quasi-Newton approach byrd2016stochastic ; berahas2016multi that approximates the Hessian using the information from past and current gradients; and the Gauss-Newton method gargiani2020promise ; ren2019efficient that leverages the Jacobian of the residuals, neglecting second-order cross-derivatives, particularly suitable for loss functions structured as (3.2). In this paper, we use the Gauss-Newton Hessian approximation which is shown to provide a good approximation of the true Hessian for ML applications sagun2017empirical ; papyan2018full .
@@ 3.2 Generalized Gauss-Newton Hessian Approximation
We consider the Generalized Gauss-Newton (GGN) Hessian approximation scheme schraudolph2002fast ; bottou2018optimization ; papyan2018full which is suited for both regression and multi-class classification tasks. The GGN Hessian approximation is constructed using only first-order sensitivities (see the derivation in Appendix A) to obtain
@@
where we vertically stack individual Jacobians $\mathbf{J}_{\Phi_{i}}:=\frac{\partial\Phi(\mathbf{x}_{i};\mathbf{w})}{\partial\mathbf{w}}$ for each sample in the batch $\mathcal{B}$ to form $\mathbf{J}=\left[\begin{array}[]{ccc}\mathbf{J}_{\Phi_{1}}&\dots&\mathbf{J}_{\Phi_{b}}\end{array}\right]^{\top}\in{\mathbb{R}}^{bc\times d}$ and construct a block diagonal matrix $\mathbf{Q}=\text{blkdiag}(\mathbf{Q}_{\ell_{1}},\mathbf{Q}_{\ell_{2}},\ldots,\mathbf{Q}_{\ell_{b}})\in{\mathbb{R}}^{bc\times bc}$, where $\mathbf{Q}_{\ell_{i}}=\frac{\partial^{2}\ell(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w}))}{\partial\Phi^{2}}\in{\mathbb{R}}^{c\times c}$. As pointed out in sankar2021deeper , approximation (3.6) is justified since the true Hessian is dominated by the term $\mathbf{J}^{\top}\mathbf{Q}\mathbf{J}$. Using the inverse of the GGN Hessian as a preconditioner yields the following direction
@@
The Gauss-Newton step (3.7) solves issues (a) and, partially, (b), since no second-order derivatives need to be computed and the Hessian approximation is positive semi-definite by construction provided that the loss function $\ell\left(\mathbf{y}_{i},\Phi(\mathbf{x}_{i};\mathbf{w})\right)$ is convex. A positive-definite Hessian approximation can easily be obtained from $\mathbf{H}^{\mathrm{GN}}$, e.g., by adding to it a small constant times the identity matrix. This approach is called Levenberg-Marquardt (LM) nocedal2006numerical and is often used in practice, such that
@@
with $\lambda>0$. The regularizer $\lambda\mathbf{I}_{d}$ serves a dual purpose: it ensures that the approximate Hessian matrix is invertible and also it helps to avoid over-fitting pml1Book .
@@
We can show that the gradient and the GN Hessian for the MSE loss read
@@
The gradient and the GN Hessian for the CE loss are
@@
where $\mathbf{J}\in{\mathbb{R}}^{bc\times d}$ is a matrix of stacked Jacobians, $\mathbf{r}\in{\mathbb{R}}^{bc}$ is a vector of (pseudo-)residuals defined as $\mathbf{r}:=\left[\begin{array}[]{ccc}\left(\sigma\left(\Phi(\mathbf{x}_{1};\mathbf{w})\right)-\mathbf{y}_{1}\right)^{\top}&\dots&\left(\sigma\left(\Phi(\mathbf{x}_{b};\mathbf{w})\right)-\mathbf{y}_{b}\right)^{\top}\end{array}\right]^{\top}$, and $\mathbf{Q}\in{\mathbb{R}}^{bc\times bc}$ is a block diagonal matrix of stacked matrices $\mathbf{Q}_{\ell_{i}}$ that each have $\sigma(\mathbf{z}_{i,k})(1-\sigma(\mathbf{z}_{i,k}))$ across the diagonal and $\sigma(\mathbf{z}_{i,k})\sigma(\mathbf{z}_{i,l})$ off-diagonal.